The cost of a trip is RM600. The cost for each member would be RM5 less if 10 non members were allowed to join. If the cost was shared equally among the students, find the number of members going for the trip
Answer by mananth(16946)

let the number of members be x
cost per person = 600/x
if 10 more join cost per person will be 600/ x+10
600/x+10 +5 = 600/x
600/x+10 - 600/x = -5
600x-600x-6000= -5x(x+10)
-6000=-5x^2-50x
5x^2+50x-6000=
x^2+10x-1200=0
x^2+40x-30x-1200=0
x(x+40)-30(x+40)=0
(x-30)(x+40)=0
x=30 the number of members